OpenStudy (anonymous):

a^2/ab^2 solve algebra

OpenStudy (mathstudent55):

If this is the problem, \( \dfrac{a^2}{ab^2} \) then use the rule: \(\dfrac{x^m}{x^n} = x^{m - n} \)